# Bringing closure to FDR control: beating the e-Benjamini-Hochberg procedure

### v1

@@ 5 Extensions
Post-hoc validity.
E-values have a notion of post-hoc validity that allows one to derive guarantees for a data-dependent choice of $\alpha$ (Grünwald, 2024; Koning, 2024), which the eBH procedure also possesses. This property is preserved for ${\overline{\text{eBH}}}$ as well.
Proposition 4.
The ${\overline{\text{eBH}}}$ procedure is valid for post-hoc (data-dependent) choice of $\alpha\in(0,1]$ in the following sense:
$\mathbb{E}\left[\frac{\textnormal{FDP}_{A^{*}}(R^{{\overline{\textnormal{eBH}}}_{\alpha}})}{\alpha}\right]\leq 1,$ (20)
where $R^{{\overline{\textnormal{eBH}}}_{\alpha}}$ is the ${\overline{\text{eBH}}}$ procedure applied at level $\alpha$.
Proof.
By construction, we have the following deterministic inequality
$\frac{\textnormal{FDP}_{A^{*}}(R^{{\overline{\textnormal{eBH}}}_{\alpha}})}{\alpha}\leq E_{A^{*}}$ (21)
that is satisfied even for possibly data-dependent $\alpha$. We get our desired result by taking expectations on both sides. ∎
@@
Recall that $R_{k}$ denotes the set of hypotheses corresponding to the $k$ largest e-values, and solve for $\hat{k}:=k^{\overline{\textnormal{eBH}}}$ from (5). Then, define data-dependent test levels
$\hat{\alpha}_{A,k}\coloneqq\left(\frac{\textnormal{FDP}_{A}(R_{k})}{\alpha}\right)^{-1}$ (22)
@@ 6 Numerical simulations
We perform simulations for independent Gaussians where we let $\pi_{0}\coloneqq|A^{*}|/K$ be the null proportion, $\mu$ be the signal of the alternatives and let $X_{i}\sim\mathcal{N}(0,1)$ for $i\in A^{*}$ and $X_{i}\sim\mathcal{N}(\mu,1)$ otherwise for $i\in[K]\setminus A^{*}$. We let our e-value be $E_{i}\coloneqq\exp(\lambda|X_{i}|-\lambda^{2}/2)/2$ for $\lambda=\sqrt{\log(2K/\alpha)}$. We let $\alpha=0.1$ in our simulations. We plot the empirical average FDR and TDP $\coloneqq\mathbb{E}[|A^{*}\cap R|/|A^{*}|]$ of eBH, minimally adaptive eBH (Ignatiadis et al., 2024), and ${\overline{\text{eBH}}}$ over $n=1000$ trials. We plot the results in Figure 1, and can see that the ${\overline{\text{eBH}}}$ procedure improves noticeably over the standard and minimally adaptive eBH procedures across all settings, while controlling the FDR below $\alpha$.
Figure 1: Plot of empirical FDR and TDP for different null proportions $\pi_{0}$ and alternative mean $\mu=3$. All differences are significant since the closed eBH ($\overline{\textnormal{eBH}}$) procedure dominates the eBH and minimally adaptive eBH (eBH min-ada) procedures.
@@ 7 Related work
There is a vast amount of literature for FWER control and simultaneous (over all potential discovery sets) FDP bounds using closed testing — see overviews presented in Cui et al. (2021, Chapters 9 and 14) and Henning and and Westfall (2015). We will point out two specific works that are particularly relevant to ours since they also utilize e-values in the context of closed testing. The first is Vovk and Wang (2023), who formulate a method for FWER control via the arithmetic mean e-merging function, and consequently also simultaneous FDP bounds. Hartog and Lei (2025) extend the benefit of using arithmetic mean e-merging functions to graphical approaches for FWER control (Bretz et al., 2009), and derive computationally efficient algorithms for e-value versions of the Fallback (Wiens and and Dmitrienko, 2005) and Holm (1979) procedures. Our work considers the FDR metric (rather than FWER or probabilistic bounds on the FDP) which has not been connected to intersection hypothesis testing or e-merging functions before.
@@
For the first area, we have already discussed the connection between our work and the stochastic rounding framework of Xu and Ramdas (2023) in Section 5. Another particularly relevant work is that of Ignatiadis et al. (2024, Section S3.2), which introduces a minimally adaptive eBH procedure where one first tests the global null with the the arithmetic mean of the e-values, and only once the global null rejects, the eBH procedure is applied at level $\alpha K/(K-1)$. They show that this method dominates eBH as well. This method is, in turn, dominated by the ${\overline{\text{eBH}}}$ procedure since rejection of the global null implies that $E_{[K]}\geq\alpha^{-1}\geq\textnormal{FDP}(R)/\alpha$ is satisfied for all $R\subseteq[K]$ (where $E_{[K]}$ is the arithmetic mean of all e-values as defined in (2)). One can then show that the $R$ which results from eBH applied at a level of $\alpha K/(K-1)$ implies that $R\in\overline{\mathcal{C}}$ using an argument similar to that in the proof of Proposition 2. We see the empirical implications of this in our simulations in Section 6.
@@ 8 Conclusion
We have shown in this manuscript that the eBH procedure is inadmissible, and can be improved through a FDR closure principle that results in the ${\overline{\text{eBH}}}$ procedure. We provide a computationally efficient way of computing the ${\overline{\text{eBH}}}$ discovery set. Further, our FDR closure principle dictates that all FDR controlling procedures rely on e-values for testing each possible intersection hypothesis of the base hypotheses. These results open many lines of future work, particularly in practical and computationally efficient applications of the closure principle to improving existing FDR controlling procedures.

### v2

@@ 5 Extensions
Recall that $R_{[:k]}$ denotes the set of hypotheses corresponding to the $k$ largest e-values, and solve for $\hat{k}:=k^{\overline{\textnormal{eBH}}}$ from (5). Then, define data-dependent test levels
$\hat{\alpha}_{A,k}\coloneqq\left(\frac{\textnormal{FDP}_{A}(R_{[:k]})}{\alpha}\right)^{-1}$ (42)
@@ 6 Numerical simulations
Figure 1: Plot of FDR and TPR for different null proportions $\pi_{0}$ and alternative mean $\mu=3$. All differences are significant since the closed eBH ($\overline{\textnormal{eBH}}$) procedure dominates the eBH and minimally adaptive eBH (eBHm) procedures.
We perform simulations [^1] Code at https://github.com/neilzxu/closed-ebh for independent Gaussians where we let $\pi_{0}\coloneqq|A^{*}|/K$ be the null proportion, $\mu$ be the signal of the alternatives and let $X_{i}\sim\mathcal{N}(0,1)$ for $i\in A^{*}$ and $X_{i}\sim\mathcal{N}(\mu,1)$ otherwise for $i\in[K]\setminus A^{*}$. We let our e-value be $E_{i}\coloneqq\exp(\lambda X_{i}-\lambda^{2}/2)$ for $\lambda=\mu$ (the log-optimal choice). We let $\alpha=0.1$ in our simulations. We plot the empirical average FDR and true positive rate (TPR) $\coloneqq\mathbb{E}[|(A^{*})^{c}\cap R|/|(A^{*})^{c}|]$ of eBH, minimally adaptive eBH (Ignatiadis et al., 2024), and ${\overline{\text{eBH}}}$ over $n=1000$ trials. We plot the results in Figure 1, and can see that the ${\overline{\text{eBH}}}$ procedure improves noticeably over the standard and minimally adaptive eBH procedures across all settings, while controlling the FDR below $\alpha$.
Figure 2: Plot of FDR and TPR for different null proportions $\pi_{0}$ and alternative mean $\mu=3$. All differences are significant since the closed BY ($\overline{\text{BY}}$) procedure dominates other variants of the BY procedure.
We also evaluate the BY procedure, the application of minimally adaptive eBH to the e-values defined in (2.1), and the $\overline{\text{BY}}$ procedure where we set $P_{i}\coloneqq 1-\Phi(X_{i})$ and $\Phi$ is the c.d.f. of the standard normal distribution. Here, we draw $(X_{1},\dots,X_{K})$ from a multivariate normal with the same component means as described above and with Toeplitz-like covariance matrix where $|\text{Cov}(X_{i},X_{j})|=\exp(-|i-j|/10)/5$ with the covariance being positive if $|i-j|$ is even and negative otherwise. Similarly, our results in Figure 2 show that $\overline{\text{BY}}$ has substantial power increase over the other BY procedures.
@@ 7 Related work
There is a vast amount of literature for FWER control and simultaneous (over all potential discovery sets) FDP bounds using closed testing — see overviews presented in Cui et al. (2021, Chapters 9 and 14) and Henning and and Westfall (2015). We will point out two specific works that are particularly relevant to ours since they also utilize e-values in the context of closed testing. The first is Vovk and Wang (2023), who formulate a method for FWER control via the arithmetic mean e-merging function, and consequently also simultaneous FDP bounds. Hartog and Lei (2025) extend the benefit of using arithmetic mean e-merging functions to graphical approaches for FWER control (Bretz et al., 2009), and derive computationally efficient algorithms for e-value versions of the Fallback (Wiens and and Dmitrienko, 2005) and Holm (1979) procedures. Our work proposes new procedures the FDR metric (rather than FWER or probabilistic bounds on the FDP) which has not been connected to intersection hypothesis testing or e-merging functions before, and we show existing procedures (such as e-Holm) can be formualted using the generalized closure principle we develop in Section 4.
@@
For the first area, we have already discussed the connection between our work and the stochastic rounding framework of Xu and Ramdas (2023) in Section 5, as well as the minimally adaptive eBH procedure of Ignatiadis et al. (2024) in Section 2.
@@ 8 Conclusion
We have shown in this manuscript that the eBH procedure is inadmissible, and can be improved through a FDR closure principle that results in the ${\overline{\text{eBH}}}$ procedure. We provide a computationally efficient way of computing the ${\overline{\text{eBH}}}$ discovery set. Further, our FDR closure principle dictates that all FDR controlling procedures rely on e-values for testing each possible intersection hypothesis of the base hypotheses. Moreover, we generalize our FDR closure principle to general error metrics, which particularly allows to choose the error metric post-hoc. These results open many lines of future work .